OpenStudy (anonymous):

What vector is the projection of u=(1;3) on v=(4;4)? The answer is supposed to be in the coordinates. I got no clue how to go along. :/

OpenStudy (anonymous):

Except for the theorem u_v=(u*v)/(v^2)*v But I do not know how to properly use it.

OpenStudy (amistre64):



OpenStudy (amistre64):

v times |u| cos(a)

OpenStudy (amistre64):

divided by that is .. not times :)

OpenStudy (amistre64):

cos(a) = u.v/|u||v|

OpenStudy (amistre64):

u=(1;3) |u|=sqrt(10) v=(4;4) |v| = sqrt(32) u.v = 4+12 = 16 cos(a) = 16/sqrt(320)

OpenStudy (amistre64):

but then: |u|*u.v ----- = u.v/|v| |u||v| v divided by 16/sqrt(32)

OpenStudy (anonymous):

\[\frac{\overrightarrow{v}\cdot \overrightarrow{u}}{|\overrightarrow{u}|^2}\times \overrightarrow{u}\] damn that took a long time to write !!

OpenStudy (anonymous):

Would that be a matrix? ;o

OpenStudy (amistre64):

<4,4> ---------- and simplify 16/sqrt(32)

OpenStudy (anonymous):

I feel so embarassingly stupid regarding vector geometry. :C

OpenStudy (anonymous):

<4/(16/sqrt32);4/(16/sqrt32)> or is this not how you're supposed to do it?

OpenStudy (amistre64):

i think its good, just simplify the results

OpenStudy (anonymous):

i think what i wrote was the projection of v on to u lets try it. \[\overrightarrow{u}\cdot\overrightarrow{v}=1\times 4+3\times 4=16\] \[|\overrightarrow{u}|^2=32\] hmm i am getting a different answer. maybe i am wrong

OpenStudy (anonymous):

But that fetches me <sqrt2;sqrt2> right? The answer is supposed to be <2;2>

OpenStudy (experimentx):

looks like it must have magnitude of U.V and parallel to V

OpenStudy (anonymous):

that is what i get yes.

OpenStudy (anonymous):

\[\frac{\overrightarrow{u}\cdot\overrightarrow{v}}{|u^2|}=\frac{1\times 4+3\times 4}{4^2+4^2}=\frac{16}{32}=\frac{1}{2}\]

OpenStudy (amistre64):

i dint unit length the v; thats prolly an error in mine

OpenStudy (anonymous):

\[\frac{1}{2}<4,4>=<2,2>\]

OpenStudy (experimentx):

\[\sqrt{x^2 + y ^2} = 16 \] and \[ y/x = 1\]

OpenStudy (anonymous):

this be the formula \[\frac{\overrightarrow{v}\cdot \overrightarrow{u}}{|\overrightarrow{u}|^2}\times \overrightarrow{u}\] as i have seen it written

OpenStudy (amistre64):

u on v tho would be v/|v| not u/|u| right?

OpenStudy (anonymous):

The |v^2| really explains alot. ^^ I do not know why they're skipping the abs operator in the demoninator - in the theorem.

OpenStudy (amistre64):

\[|u|\frac{u.v}{|u||v|}*\frac{v}{|v|}\]

OpenStudy (anonymous):

Superb. A thousand thanks! :)

OpenStudy (experimentx):

I've never seen that formula ... what exactly is it??